Corporation A has $81940 in taxable income, and Corporation B has $7.9 million in taxable income. Using the tax rates from Table 2.3, what is the difference between the tax bills of these two firms?

(Do not round intermediate calculations and round your final answer to the nearest dollar amount. Omit the "$" sign and commas in your response. For example, $123,456.78 should be entered as 123457.)

TABLE 2.3 Taxable Income Tax Rate Corporate Tax Rates 50,000 75,000 100,000 100,001335,000 335,001-10,000,000 10,000,001-15,000,000 15,000,001-18,333,333 15% 25 34 39 34 35 38 35 50,001 75,001 18,333,334+
Answer #1

Answer:

Corporation A has $81940 in taxable income

Tax bill of corporate A = 15% * 50000 + 25% * (75000 - 50000) + 34% * (81940 - 75000) = 16109.6

Tax bill of corporate B:

= 15% * 50000 + 25% * (75000 - 50000) + 34% * (100000 - 75000) + 39% * (335000 - 100000) + 34% * (7900000 - 335000)

=2686000

Difference between the tax bills of these two firms = 2686000 - 16109.6 = 2669890

Difference between the tax bills of these two firms = 2669890